OpenStudy (anonymous):

simplify (x^3-8)/(x^2-4)

OpenStudy (anonymous):

x^3-a^3=(x-a)(x^2+xa+a^2) x^2-a^2=(x-a)(x+a)

OpenStudy (anonymous):

wait...where does the A come from ?

OpenStudy (anonymous):

it's formula. U replace "a" with number in Ur problem

OpenStudy (anonymous):

U can rewrite 8 as something in 3rd power

OpenStudy (anonymous):

I don't know what this is. sorry, i still don't understand.

OpenStudy (anonymous):

8=2^3, replace a^3 with 2^3